OpenStudy (anonymous):

A volume of 90.0mL of a 0.200M HNO3 solution is titrated with 0.800M KOH. Calculate the volume of KOH required to reach the equivalence point.

OpenStudy (aaronq):

the moles of \(H^+\) should equal the moles of \(OH^-\). Each of these releases 1 equivalent of their corresponding ions, therefore you can use: \(M_1V_1=M_2V_2\) ..if you don't see it, then write an equation for the neutralization reaction and find the moles of KOH equivalent to the moles of \(HNO_3\) . Use a ratio for this: Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) once you find the moles of KOH you can find the volume. \(M=\dfrac{n_{solute}}{L_{solution}}\)